OpenStudy (anonymous):

What is 3 root of 125 ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

thank you can you explain why?

OpenStudy (anonymous):

\[5*5*5 = 5^{3}\]

OpenStudy (anonymous):

=125

OpenStudy (anonymous):

thank you ^_^

OpenStudy (igreen):

That's the cube root.

OpenStudy (butterflydreamer):

you can plug it into a calculator :) Or remember that 125 = 25 * 5 = 5 * 5 * 5 cube root means they want to find a number CUBED that equals to 125 :)

OpenStudy (igreen):

It's asking, what number multiplied by itself 3 times equals 125.

OpenStudy (igreen):

It's kind of like the square root..

OpenStudy (igreen):

@Ziferot Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

@iGreen ok, i'll do my best)